OpenStudy (anonymous):

the length of the hypotensue of a 30 -60 -90 triangle is 28m. Find the length of the side opposite the 30degree angle

Parth (parthkohli):

28/3 = 9.33 , not sure... still

OpenStudy (anonymous):

ratios are \(1:\sqrt{3}:2\) for \(\text{short side}: \text{long side}:\text{hypotenuse}\)

OpenStudy (anonymous):

I will be back on late

OpenStudy (anonymous):

you want short side so \(\frac{28}{2}\)

OpenStudy (anonymous):

a..7m b.14m c.14 3 squaredm d.14 2 squared meter which one is it?

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

ok thanks